# Accelerating Quantum Reinforcement Learning with a Quantum Natural Policy Gradient Based Approach

### v1

@@ 1 Introduction
Reinforcement Learning (RL) is a foundational framework for sequential decision-making, with applications spanning robotics, finance, transportation, and healthcare (Al-Abbasi et al., 2019; Gonzalez et al., 2023; Tamboli et al., 2024; Wang et al., 2023). The goal of an RL agent is to derive a policy that maximizes the discounted sum of expected rewards through interactions with the environment. A prominent approach to solving this problem is the policy gradient (PG) method, which optimizes directly in the policy space. In classical RL, the optimal sample complexity of this approach has been established as $\tilde{\mathcal{O}}(\epsilon^{-2})$ (Mondal and Aggarwal, 2024). Quantum computing has shown potential for speedups in areas such as mean estimation (Hamoudi, 2021), multi-armed bandits (Wan et al., 2023; Wu et al., 2023), and tabular reinforcement learning (Ganguly et al., 2024; Zhong et al., 2024). This work aims to explore potential quantum speedups for reinforcement learning with general parameterized policies.
@@
Quantum computing principles have shown significant improvements in areas such as quantum mean estimation (Hamoudi, 2021) and quantum stochastic convex optimization (Sidford and Zhang, 2024). However, quantum reinforcement learning (RL) presents unique challenges. Unlike classical algorithms, quantum methods rely on deterministic operations, complicating the design of unbiased estimators for policy gradients. This paper addresses the trade-off between bias and unbiased estimation by introducing a novel deterministic algorithm that facilitates quantum-compatible gradient estimation. While this approach introduces a bounded bias, it achieves orders reduction in sample complexity compared to its classical counterparts.
@@ 1.1 Challenges and Contributions
In this work, we introduce the first quantum model-free reinforcement learning (RL) algorithm with theoretical guarantees, addressing the challenges posed by large state and action spaces. To our knowledge, this is the first work to coherently embed the entire Natural Policy Gradient (NPG) into a quantum state by leveraging only the standard environment oracles from reinforcement learning. This novel construction allows subsequent quantum subroutines to utilize these NPG gradients directly in superposition, enabling potential accelerations in policy optimization. Extending classical model-free RL algorithms to the quantum domain is nontrivial. Specifically, classical model-free algorithms (Mondal and Aggarwal, 2024; Liu et al., 2020; Agarwal et al., 2021) typically rely on sampling trajectories of random lengths following a geometric distribution for policy gradient estimation. However, encoding such trajectories with random lengths into quantum systems is challenging. To address this, we propose a novel deterministic sampling algorithm that uses truncation to estimate both the Fisher matrix and the policy gradient. We further develop a mini-batch strategy to incorporate quantum variance reduction into the Natural Policy Gradient (NPG) update. By carefully analyzing the bias introduced by the truncation, we show that our approach achieves a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-1.5})$, surpassing the classical lower bound of $\tilde{\mathcal{O}}(\epsilon^{-2})$. To the best of our knowledge, this is the first work to demonstrate quantum speedups for parameterized model-free infinite-horizon Markov Decision Processes (MDPs).
It should be noted that SGD-based approaches have been shown to achieve a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ for stochastic convex optimization (Sidford and Zhang, 2024). However, reinforcement learning (RL) presents a fundamentally different set of challenges compared to stochastic convex optimization. While insights from the literature suggest that $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ is likely the best achievable sample complexity for SGD-based approaches, this limitation arises because quantum computing can accelerate the inner loop but not the outer loop of such algorithms. To achieve better sample complexity guarantees, stochastic cutting-plane methods have been proposed in (Sidford and Zhang, 2024). Extending these methods to RL, however, would require an entirely new algorithmic framework. This is due to the non-convex nature of RL with general parameterization, which makes direct adaptation infeasible. Moreover, the techniques used in (Agarwal et al., 2021) to transition from local to global convergence are not applicable in this context, preventing similar guarantees for such methods in RL.
@@ 1.3 Related Work and Preliminaries
The unitary operator $\mathcal{U}_{\mathbb{P}}$ enables the query of samples of the random variable in superposition, forming the basis for the speed-ups achieved in quantum algorithms. To perform quantum mean estimation for random variable values Cornelissen et al. (2022); Hamoudi (2021); Montanaro (2015), the implementation of an additional quantum evaluation oracle is required.
@@
Quantum Reinforcement Learning: In recent years, Quantum Reinforcement Learning (QRL) has garnered substantial interest from the research community (Jerbi et al., 2021; Dunjko et al., 2017; Paparo et al., 2014). Within this field, studies on the Quantum Multi-Armed Bandits (Q-MAB) problem have demonstrated exponential reductions in regret by leveraging amplitude amplification techniques (Wang et al., 2021b; Casalé et al., 2020). However, the theoretical foundations of Q-MAB approaches are not directly applicable to QRL, primarily due to the lack of state transitions in bandit problems.
Although interest in QRL has been growing, most existing works are still limited to the tabular setup. For instance, (Zhong et al., 2024; Ganguly et al., 2024) showed that logarithmic regret is achievable in episodic/average reward Markov Decision Processes (MDPs) using model-based approaches, while Wang et al. (2021a) investigated the infinite horizon discounted MDP framework with tabular model-free setups. However, the results don’t directly extend to large state spaces, where a parameterized policy is used. Further, the problem becomes non-convex due to the use of general parameterization. To the best of our knowledge, ours is the first work to address infinite horizon MDPs with general parameterized policies.
@@ 2.1 Quantum Computing Basics:
In this section, we provide a concise overview of the key concepts relevant to our work, drawing from (Nielsen and Chuang, 2010) and (Wu et al., 2023). Let $\mathbb{C}^{m}$ denote an $m$-dimensional Hilbert space. A quantum state $|x\rangle=(x_{1},\ldots,x_{m})^{T}$ can be represented as a vector within this space, satisfying the normalization condition $\sum_{i}|x_{i}|^{2}=1$. For a finite set $\xi=\{\xi_{1},\ldots,\xi_{m}\}$ of $m$ elements, each $\xi_{i}\in\xi$ can be mapped to an element of an orthonormal basis in $\mathbb{C}^{m}$ as follows:
@@ 2.2 MDP Formulation
$Q^{\pi}(s,a)\triangleq\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\bigg{|}s_{0}=s,a_{0}=a\right]$
@@
$V^{\pi}(s)\triangleq\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\bigg{|}s_{0}=s\right]=\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}(s,a)$
@@
$J^{\pi}_{\rho}\triangleq\mathbb{E}_{s\sim\rho}\left[V^{\pi}(s)\right]=\dfrac{1}{1-\gamma}\sum_{s,a}d^{\pi}_{\rho}(s,a)\pi(a|s)r(s,a)$
@@ 2.3 Quantum access to an MDP
We adopt the framework and notations of quantum computing from Zhong et al. (2024); Wiedemann et al. (2022); Ganguly et al. (2024); Jerbi et al. (2023) to design the quantum sampling oracle for reinforcement learning (RL) environments, enabling the modeling of an agent’s interactions with an unknown MDP environment. We proceed to define quantum-accessible RL environments corresponding to the classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and performing action $a_{t}$, we construct quantum sampling oracles for the transition probabilities of the next state, $P(\cdot|s_{t},a_{t})$. To this end, let two Hilbert spaces, $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$, represent the superpositions of classical states and actions, respectively. The computational bases for these Hilbert spaces are denoted as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume the ability to implement the following quantum sampling oracles:
@@
This construction is particularly intuitive for certain quantum policies, such as raw-PQC which is introduced in Jerbi et al. (2023). However, any policy that can be classically computed can also be converted into such a unitary operation through quantum simulation of the classical computation of $(\pi_{\theta}(a|s):a\in\mathcal{A})$ and leveraging established methods to encode this probability vector into the amplitudes of a quantum state (Grover and Rudolph, 2002). With access to quantum oracles for both the environment and the policy in (4)-(6), it becomes possible to design subroutines capable of generating superpositions of trajectories with fixed length within the environment and calculating the returns associated with these trajectories. We can further utilize $\mathcal{U}_{P}$, $\mathcal{U}_{\rho}$ and $\Pi$ defined above to construct a unitary $\mathcal{U}_{P(\tau_{N})}$ as follows: Let $\mathcal{M}$ be a quantum-accessible MDP with oracles $\mathcal{U}_{P},\mathcal{U}_{\rho}$ as above, and let $\pi_{\theta}$ be a quantum-evaluatable policy with its unitary implementation $\Pi$ as defined above. Given a fixed number $N$, the unitary $\mathcal{U}_{P(\tau_{N})}$ that prepares a coherent superposition of all trajectories $\tau_{N}=(s_{0},a_{0},\ldots,s_{N-1},a_{N-1})$ of length $N$ (without their rewards) is defined as ,
@@ 3 Proposed Algorithm
$\nabla_{\theta}J_{\rho}(\theta)=\dfrac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim\nu_{\rho}^{\pi_{\theta}}}\big{[}A^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\big{]}$ (8)
@@
$\begin{split}F_{\rho}(\theta)\triangleq\mathbb{E}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\big{[}\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\big{]}\end{split}$ (10)
@@
2:for $k\in\{0,\cdots,K-1\}$ do $\triangleright$ Outer Loop
3:     $\mathbf{x}_{0},\mathbf{v}_{0}\leftarrow\mathbf{0}$
4:     for $h\in\{0,\cdots,H-1\}$ do $\triangleright$ Inner Loop
5:         $\triangleright$ Mini-Batch Quantum Stochastic Gradient Descent
6:         Obtain gradient estimator $\tilde{g}_{h}\leftarrow\texttt{QVarianceReduce}(\hat{g}_{\rho}(\tau_{N}|\theta_{k}),\hat{\sigma}^{2}_{g})$
7:         Obtain Fisher estimator $\tilde{F}_{h}\leftarrow\texttt{QVarianceReduce}(\hat{F}_{\rho}(\tau_{N}|\theta_{k}),\hat{\sigma}^{2}_{F})$
8:         $\nabla_{\omega}\tilde{L}(\omega,\tau_{N})\leftarrow\tilde{F}_{h}\omega-\tilde{g}_{h}$
9:         NPG update: $\omega_{h+1}=\omega_{h}-\alpha\nabla_{\omega}\tilde{L}(\omega,\tau_{N})\big{|}_{\omega=\omega_{h}}$
@@
$\begin{split}~{}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta)\triangleq\dfrac{1}{2}\mathbb{E}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\bigg{[}\dfrac{1}{1-\gamma}A^{\pi_{\theta}}(s,a)-\omega^{\mathrm{T}}\nabla_{\theta}\log\pi_{\theta}(a|s)\bigg{]}^{2}\end{split}$ (12)
@@
We now present an estimation procedure for obtaining quantum estimators for $\nabla_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta)$ based on the provided oracles in Section 2.3. Different from the commonly used unbiased estimators in (Agarwal et al., 2021; Mondal and Aggarwal, 2024) which requires sampling from geometric distribution with mean $\frac{1}{1-\gamma}$, we construct a deterministic algorithm which can be implemented using the quantum oracles in Section 2.3. For a fixed $N$, let $\tau_{N}=(s_{0},a_{0},\ldots,s_{N},a_{N})$ be a trajectory with length $N$ generated by taking policy $\pi_{\theta}$ with $s_{0}$ sampled from $\rho$, we construct the policy gradient estimator $\hat{g}_{\rho}(\tau_{N}|\theta)$ and Fisher information matrix estimator $\hat{F}_{\rho}(\tau_{N}|\theta)$ as follows:
@@
$\begin{split}\hat{F}_{\rho}(\tau_{N}|\theta)=(1-\gamma)\sum_{n=0}^{N-1}\gamma^{n}\Bigl{(}\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\otimes\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\Bigr{)},\end{split}$ (15)
@@
$\mathcal{U}_{F}\left(|\Psi_{\theta}\rangle\otimes|0\rangle\right)=|\psi_{F}^{\theta}\rangle\coloneqq|\theta\rangle\sum_{\tau_{N}}\sqrt{P_{\theta}(\tau_{N})}|\tau_{N}\rangle\otimes\left|\hat{F}_{\rho}(\tau_{N}|\theta)\right\rangle,$ (16)
$\mathcal{U}_{g}\left(|\Psi_{\theta}\rangle\otimes|0\rangle\right)=|\psi_{g}^{\theta}\rangle\coloneqq|\theta\rangle\sum_{\tau_{N}}\sqrt{P_{\theta}(\tau_{N})}|\tau_{N}\rangle\otimes\left|\hat{g}_{\rho}(\tau_{N}|\theta)\right\rangle,$ (17)

### v3

@@ 1 Introduction
Reinforcement Learning (RL) is a foundational framework for sequential decision-making, with applications spanning robotics, finance, transportation, and healthcare (Al-Abbasi et al., 2019; Gonzalez et al., 2023; Tamboli et al., 2024; Wang et al., 2023). The goal of an RL agent is to derive a policy that maximizes the discounted sum of expected rewards through interactions with the environment. A prominent approach to solving this problem is the policy gradient (PG) method, which optimizes directly in the policy space. In classical RL, the optimal sample complexity of this approach has been established as $\tilde{\mathcal{O}}(\epsilon^{-2})$ (Mondal & Aggarwal, 2024a). Quantum computing has shown potential for speedups in areas such as mean estimation (Hamoudi, 2021), multi-armed bandits (Wan et al., 2023; Wu et al., 2023), and tabular reinforcement learning (Ganguly et al., 2023; Zhong et al., 2024). This work aims to explore potential quantum speedups for reinforcement learning with general parameterized policies.
@@
Quantum computing principles have shown significant improvements in areas such as quantum mean estimation (Hamoudi, 2021) and quantum stochastic convex optimization (Sidford & Zhang, 2024). However, quantum reinforcement learning (RL) presents unique challenges. Unlike classical algorithms, quantum methods rely on deterministic operations, complicating the design of unbiased estimators for policy gradients. This paper addresses the trade-off between bias and unbiased estimation by introducing a novel deterministic algorithm that facilitates quantum-compatible gradient estimation. While this approach introduces a bounded bias, it achieves orders reduction in sample complexity compared to its classical counterparts.
@@ 1.1 Challenges and Contributions
In this work, we introduce the first quantum model-free reinforcement learning (RL) algorithm with theoretical guarantees, addressing the challenges posed by large state and action spaces. To our knowledge, this is the first work to coherently embed the entire Natural Policy Gradient (NPG) into a quantum state by leveraging only the standard environment oracles from reinforcement learning. This novel construction allows subsequent quantum subroutines to utilize these NPG gradients directly in superposition, enabling potential accelerations in policy optimization. Extending classical model-free RL algorithms to the quantum domain is nontrivial. Specifically, classical model-free algorithms (Mondal & Aggarwal, 2024a; Liu et al., 2020; Agarwal et al., 2021) typically rely on sampling trajectories of random lengths following a geometric distribution for policy gradient estimation. However, encoding such trajectories with random lengths into quantum systems is challenging. To address this, we propose a novel deterministic sampling algorithm that uses truncation to estimate both the Fisher matrix and the policy gradient. We further develop a mini-batch strategy to incorporate quantum variance reduction into the Natural Policy Gradient (NPG) update. By carefully analyzing the bias introduced by the truncation, we show that our approach achieves a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-1.5})$, surpassing the classical lower bound of $\tilde{\mathcal{O}}(\epsilon^{-2})$. To the best of our knowledge, this is the first work to demonstrate quantum speedups for parameterized model-free infinite-horizon Markov Decision Processes (MDPs).
It should be noted that SGD-based approaches have been shown to achieve a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ for stochastic convex optimization (Sidford & Zhang, 2024). However, reinforcement learning (RL) presents a fundamentally different set of challenges compared to stochastic convex optimization. While insights from the literature suggest that $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ is likely the best achievable sample complexity for SGD-based approaches, this limitation arises because quantum computing can accelerate the inner loop but not the outer loop of such algorithms. To achieve better sample complexity guarantees, stochastic cutting-plane methods have been proposed in (Sidford & Zhang, 2024). Extending these methods to RL, however, would require an entirely new algorithmic framework. This is due to the non-convex nature of RL with general parameterization, which makes direct adaptation infeasible. Moreover, the techniques used in (Agarwal et al., 2021) to transition from local to global convergence are not applicable in this context, preventing similar guarantees for such methods in RL.
@@ 1.3 Related Work and Preliminaries
The unitary operator $\mathcal{U}_{\mathbb{P}}$ enables the query of samples of the random variable in superposition, forming the basis for the speed-ups achieved in quantum algorithms. To perform quantum mean estimation for random variable values (Cornelissen et al., 2022; Hamoudi, 2021; Montanaro, 2015), the implementation of an additional quantum evaluation oracle is required.
@@
Quantum Reinforcement Learning: In recent years, Quantum Reinforcement Learning (QRL) has garnered substantial interest from the research community (Dong et al., 2008; Paparo et al., 2014; Dunjko et al., 2017; Jerbi et al., 2021). Within this field, studies on the Quantum Multi-Armed Bandits (Q-MAB) problem have demonstrated exponential reductions in regret by leveraging amplitude amplification techniques (Wang et al., 2021b; Casalé et al., 2020). However, the theoretical foundations of Q-MAB approaches are not directly applicable to QRL, primarily due to the lack of state transitions in bandit problems.
Although interest in QRL has been growing, most existing works are still limited to the tabular setup. For instance, (Zhong et al., 2024; Ganguly et al., 2023) showed that logarithmic regret is achievable in episodic/average reward Markov Decision Processes (MDPs) using model-based approaches, while (Wang et al., 2021a) investigated the infinite horizon discounted MDP framework with tabular model-free setups. However, the results don’t directly extend to large state spaces, where a parameterized policy is used. Further, the problem becomes non-convex due to the use of general parameterization. To the best of our knowledge, ours is the first work to address infinite horizon MDPs with general parameterized policies.
@@ 2.1 Quantum Computing Basics:
In this section, we provide a concise overview of the key concepts relevant to our work, drawing from (Nielsen & Chuang, 2010) and (Wu et al., 2023). Let $\mathbb{C}^{m}$ denote an $m$-dimensional Hilbert space. A quantum state $|x\rangle=(x_{1},\ldots,x_{m})^{T}$ can be represented as a vector within this space, satisfying the normalization condition $\sum_{i}|x_{i}|^{2}=1$. For a finite set $\xi=\{\xi_{1},\ldots,\xi_{m}\}$ of $m$ elements, each $\xi_{i}\in\xi$ can be mapped to an element of an orthonormal basis in $\mathbb{C}^{m}$ as follows:
@@ 2.2 MDP Formulation
$Q^{\pi}(s,a)\triangleq\operatorname{\mathbb{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\bigg{|}s_{0}=s,a_{0}=a\right]$
@@
$V^{\pi}(s)\triangleq\operatorname{\mathbb{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\bigg{|}s_{0}=s\right]=\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}(s,a)$
@@
$J^{\pi}_{\rho}\triangleq\operatorname{\mathbb{E}}_{s\sim\rho}\left[V^{\pi}(s)\right]=\dfrac{1}{1-\gamma}\sum_{s,a}d^{\pi}_{\rho}(s,a)\pi(a|s)r(s,a)$
@@ 2.3 Quantum access to an MDP
We adopt the framework and notations of quantum computing from (Zhong et al., 2024; Wiedemann et al., 2022; Ganguly et al., 2023; Jerbi et al., 2023) to design the quantum sampling oracle for reinforcement learning (RL) environments, enabling the modeling of an agent’s interactions with an unknown MDP environment. We proceed to define quantum-accessible RL environments corresponding to the classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and performing action $a_{t}$, we construct quantum sampling oracles for the transition probabilities of the next state, $P(\cdot|s_{t},a_{t})$. To this end, let two Hilbert spaces, $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$, represent the superpositions of classical states and actions, respectively. The computational bases for these Hilbert spaces are denoted as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume the ability to implement the following quantum sampling oracles:
@@
This construction is particularly intuitive for certain quantum policies, such as raw-PQC which is introduced in (Jerbi et al., 2023). However, any policy that can be classically computed can also be converted into such a unitary operation through quantum simulation of the classical computation of $(\pi_{\theta}(a|s):a\in\mathcal{A})$ and leveraging established methods to encode this probability vector into the amplitudes of a quantum state (Grover & Rudolph, 2002). With access to quantum oracles for both the environment and the policy in (4)-(6), it becomes possible to design subroutines capable of generating superpositions of trajectories with fixed length within the environment and calculating the returns associated with these trajectories. We can further utilize $\mathcal{U}_{P}$, $\mathcal{U}_{\rho}$ and $\Pi$ defined above to construct a unitary $\mathcal{U}_{P(\tau_{N})}$ as follows: Let $\mathcal{M}$ be a quantum-accessible MDP with oracles $\mathcal{U}_{P},\mathcal{U}_{\rho}$ as above, and let $\pi_{\theta}$ be a quantum-evaluatable policy with its unitary implementation $\Pi$ as defined above. Given a fixed number $N$, the unitary $\mathcal{U}_{P(\tau_{N})}$ that prepares a coherent superposition of all trajectories $\tau_{N}=(s_{0},a_{0},\ldots,s_{N-1},a_{N-1})$ of length $N$ (without their rewards) is defined as ,
@@ 3 Proposed Algorithm
$\nabla_{\theta}J_{\rho}(\theta)=\dfrac{1}{1-\gamma}\operatorname{\mathbb{E}}_{(s,a)\sim\nu_{\rho}^{\pi_{\theta}}}\big{[}A^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\big{]}$ (8)
@@
$\begin{split}F_{\rho}(\theta)\triangleq\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\big{[}\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\big{]}\end{split}$ (10)
@@
2:  for $k\leftarrow\{0,\cdots,K-1\}$ do
3:     for $h\in\{0,\cdots,H-1\}$ do
4:        {Mini-Batch Quantum SGD}
5:        $\tilde{g}_{h}\leftarrow\texttt{QVarianceReduce}(\hat{g}_{\rho}(\tau_{N}|\theta_{k}),\hat{\sigma}^{2}_{g})$ Obtain gradient estimator
6:        $\tilde{F}_{h}\leftarrow\texttt{QVarianceReduce}(\hat{F}_{\rho}(\tau_{N}|\theta_{k}),\hat{\sigma}^{2}_{F})$ Obtain Fisher estimator
7:        $\nabla_{\omega}\tilde{L}(\omega,\tau_{N})\leftarrow\tilde{F}_{h}\omega-\tilde{g}_{h}$
8:        NPG update: $\omega_{h+1}=\omega_{h}-\alpha\nabla_{\omega}\tilde{L}(\omega,\tau_{N})\big{|}_{\omega=\omega_{h}}$
9:     end for
@@
$\begin{split}~{}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta)\triangleq\dfrac{1}{2}\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\bigg{[}\dfrac{1}{1-\gamma}A^{\pi_{\theta}}(s,a)\\ -\omega^{\mathrm{T}}\nabla_{\theta}\log\pi_{\theta}(a|s)\bigg{]}^{2}\end{split}$ (12)
@@
We now present an estimation procedure for obtaining quantum estimators for $\nabla_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta)$ based on the provided oracles in Section 2.3. Different from the commonly used unbiased estimators in (Agarwal et al., 2021; Mondal & Aggarwal, 2024a) which requires sampling from geometric distribution with mean $\frac{1}{1-\gamma}$, we construct a deterministic algorithm which can be implemented using the quantum oracles in Section 2.3. For a fixed $N$, let $\tau_{N}=(s_{0},a_{0},\ldots,s_{N},a_{N})$ be a trajectory with length $N$ generated by taking policy $\pi_{\theta}$ with $s_{0}$ sampled from $\rho$, we construct the policy gradient estimator $\hat{g}_{\rho}(\tau_{N}|\theta)$ and Fisher information matrix estimator $\hat{F}_{\rho}(\tau_{N}|\theta)$ as follows:
@@
$\begin{split}\hat{F}_{\rho}(\tau_{N}|\theta)=(1-\gamma)\sum_{n=0}^{N-1}\gamma^{n}\Bigl{(}\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\\ \quad\otimes\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\Bigr{)},\end{split}$ (15)
@@
$\mathcal{U}_{F}\left(|\Psi_{\theta}\rangle\otimes|0\rangle\right)$ $=|\psi_{F}^{\theta}\rangle$
$\coloneqq|\theta\rangle\sum_{\tau_{N}}\sqrt{P_{\theta}(\tau_{N})}|\tau_{N}\rangle\otimes\left|\hat{F}_{\rho}(\tau_{N}|\theta)\right\rangle,$ (16)
$\mathcal{U}_{g}\left(|\Psi_{\theta}\rangle\otimes|0\rangle\right)$ $=|\psi_{g}^{\theta}\rangle$
$\coloneqq|\theta\rangle\sum_{\tau_{N}}\sqrt{P_{\theta}(\tau_{N})}|\tau_{N}\rangle\otimes\left|\hat{g}_{\rho}(\tau_{N}|\theta)\right\rangle,$ (17)